# Sparse identification of nonlinear dynamics and Koopman operators with Shallow Recurrent Decoder Networks

### v1

@@ Abstract
Spatiotemporal modeling of real-world data poses a challenging problem due to inherent high dimensionality, measurement noise, and expensive data collection procedures. In this paper, we present Sparse Identification of Nonlinear Dynamics with SHallow REcurrent Decoder networks (SINDy-SHRED), a method to jointly solve the sensing and model identification problems with simple implementation, efficient computation, and robust performance. SINDy-SHRED uses Gated Recurrent Units (GRUs) to model the temporal sequence of sensor measurements along with a shallow decoder network to reconstruct the full spatiotemporal field from the latent state space using only a few available sensors. Our proposed algorithm introduces a SINDy-based regularization; beginning with an arbitrary latent state space, the dynamics of the latent space progressively converges to a SINDy-class functional, provided the projection remains within the set. In restricting SINDy to a linear model, the architecture produces a Koopman-SHRED model which enforces a linear latent space dynamics. We conduct a systematic experimental study including synthetic PDE data, real-world sensor measurements for sea surface temperature, and direct video data. With no explicit encoder, SINDy-SHRED and Koopman-SHRED enable efficient training with minimal hyperparameter tuning and laptop-level computing; further, it demonstrates robust generalization in a variety of applications with minimal to no hyperparameter adjustments. Finally, the interpretable SINDy and Koopman models of latent state dynamics enables accurate long-term video predictions, achieving state-of-the-art performance and outperforming all baseline methods considered, including Convolutional LSTM, PredRNN, ResNet, and SimVP.
@@ 1 Introduction
Modeling unknown physics is an exceptionally challenging task that is complicated further by the computational burden of high-dimensional state spaces and expensive data collection. Partial differential equations (PDEs) derived from first principles remain the most ubiquitous class of models to describe physical phenomena. However, we frequently find that the simplifying assumptions necessary to construct a PDE model can render it ineffectual for real data where the physics is multi-scale in nature, only partially known, or where first principles models currently do not exist. In such cases, machine learning (ML) methods offer an attractive alternative for learning both the physics and coordinates necessary to quantify observed spatiotemporal phenomena. Many recent efforts utilizing ML techniques seek to relax the computational burden for PDE simulation by learning surrogate models to forward-simulate or predict spatiotemporal systems. However, this new machine learning paradigm frequently exhibits instabilities during the training process, unstable roll outs when modeling future state predictions, and often yields minimal computational speedups.
@@
Shallow Recurrent Decoder networks (SHRED) [52] are a recently introduced architecture that utilize data from sparse sensors to reconstruct and predict the entire spatiotemporal domain. Similar to Takens’ embedding theorem, SHRED models trade spatial information at a single time point for a trajectory of sensor measurements across time. Previous work has shown SHRED can achieve excellent performance in examples ranging from weather forecasting, atmospheric ozone concentration modeling, and turbulent flow reconstructions.
@@
Figure 1: Illustration of the SINDy-SHRED and Koopman-SHRED architecture. SINDy-SHRED transfers the original sparse sensor signal (red) to an interpretable latent representation (purple) that falls into the SINDy-class functional. This framework can be adapted into Koopman-SHRED by restricting the library $\Theta(\cdot)$ to be linear. The shallow decoder performs a reconstruction in the pixel space.
@@
In this paper, we introduce Sparse Identification of Nonlinear Dynamics with SHallow REcurrent Decoder networks (SINDy-SHRED). SINDy-SHRED exploits the latent space of recurrent neural networks for sparse sensor modeling, and enforces interprebility via a SINDy-based functional class. In this way, SINDy-SHRED enables a robust and sample-efficient joint discovery of the governing equation and coordinate system. With the correct governing equation, SINDy-SHRED can perform an accurate long-term prediction in a learned latent space, and in turn allow for long-term forecasting in the pixel space. In restricting SINDy to a linear model, a Koopman approximation can be constructed to produce a Koopman-SHRED architecture. For practical applications, SINDy-SHRED and Koopman-SHRED are lightweight models which can perform low-rank recovery with only a few (e.g. three) active sensors, which is critical for large-scale scientific data modeling and real-time control. It does not require large amounts of data during training, thereby avoiding a common pitfall in existing ML techniques for accelerating physics simulations. SINDy-SHRED and Koopman-SHRED also exhibit remarkable training speed, even when executed on a single laptop. Furthermore, SINDy-SHRED/Koopman-SHRED are highly reproducible with minimal effort in hyperparameter tuning. The recommended network structure, hyperparameter, and training setting can generalize to many different datasets  [^1] Code implementation: https://github.com/gaoliyao/sindy-shred. In short, we demonstrate SINDy-SHRED/Koopman-SHRED to be very robust and highly applicable in many modern scientific modeling problems. In what follows, we will refer generally to the proposed architecture as SINDy-SHRED, with Koopman-SHRED as a special case.
Existing work seeking to perform data-driven, long-term forecasts of spatio-temporal phenomena typically suffers from (i) instabilities and (ii) massive computational requirements. We find that SINDy-SHRED ameliorates many of these issues because (i) it is based on a stable equation discovery and (ii) the learned model is an ODE in a learned latent space, rendering simulation more computationally efficient. We further conjecture that the fact SINDy-SHRED does not include a spatial encoder contributes to the architecture’s robustness, rendering it more difficult for the model to overfit during training.
@@
We perform a wide range of studies to demonstrate the effectiveness of SINDy-SHRED. We first apply the model on the sea surface temperature data, which is a complex real-world problem. We also consider data from a complex simulation of atmospheric chemistry, video data of flow over a cylinder, and video data of a pendulum. The ability of SINDy-SHRED to perform well on video data is an important result for so-called “GoPro physics.” With extremely small sample size and noisy environments, SINDy-SHRED achieves governing equation identification with stable long-term predictions. Finally, we demonstrate the performance of SINDy-SHRED on a chaotic 2D Kolmogorov flow (in Appendix C), finding a reasonable model even in the presenece of chaos. The contribution of our paper is three-fold.
@@
  • We propose SINDy-SHRED to incorporate symbolic understanding of the latent space of recurrent models of spatio-temporal dynamics.
  • We further analyze the latent space of case studies and propose scientific models for these systems.
  • We systematically study SINDy-SHRED and compare to popular deep learning algorithms in spatio-temporal prediction.
@@ 2 Related works
Traditionally, spatio-temporal physical phenomena are modeled by Partial Differential Equations (PDEs). To accelerate PDE simulations, recent efforts have leveraged neural networks to model physics. By explicitly assuming the underlying PDE, physics-informed neural networks [42] utilize the PDE structure as a constraint for small sample learning. However, assuming the exact form of governing PDE for real data can be a strong limitation. There have been many recent works on learning and predicting PDEs directly using neural networks [26; 31; 24; 34; 32]. Meanwhile, PDE-find [46; 37; 16] offers a data-driven approach to identify PDEs from the spatial-temporal domain. Still, the high-dimensionality and required high data quality can be prohibitive for practical applications.
In parallel, previous efforts in the discovery of physical law through dimensionality reduction techniques [8; 35; 36] provide yet another perspective on the modeling of scientific data. The discovery of physics from a learned latent space has previously been explored by [18; 11; 15; 12; 55; 30; 40], yet none of these methods consider a regularization on the latent space with no explicit encoder. Yu et al. proposed the idea of physics-guided learning [57] which combines physics simulations and neural network approximations. Directly modeling physics from video is also the subject of much research in the field of robotics [17; 50; 47], computer vision [56; 53] and computer graphics [25; 33; 54; 39], since many fields of research require better physics models for simulation and control. From the deep learning side, combining the structure of differential equations into neural networks [23; 9] has been remarkably successful in a wide range of tasks. When spatial-temporal modeling is framed as a video prediction problem, He et al. found [22] that random masking can be an efficient spatio-temporal learner, and deep neural networks can provide very good predictions for the next 10 to 20 frames [48; 51; 20; 21]. Generative models have also been found to be useful for scientific data modeling [38; 49; 6].
@@ 3 Methods
The shallow recurrent decoder network (SHRED) is a computational technique that utilizes recurrent neural networks to predict the spatial domain. [52]. The method functions by trading high-fidelity spatial information for trajectories of sparse sensor measurements at given spatial locations. Mathematically, consider a high-dimensional data series $\{{\bm{X}}_{i}\}_{i=1}^{T}\in{}^{(W\times H)\otimes T}$ that represents the evolution of a spatio-temporal dynamical system, $W$, $H$, and $T$ denote the width, height, and total time steps of the system, respectively. In SHRED, each sensor collects data from a fixed spatial position in a discretized time domain. Denote the subset of sensors as $\mathcal{S}$, the input data of SHRED is $\{{\bm{X}}^{\mathcal{S}}\}_{i=1}^{T}\in{}^{\text{card}(S)\otimes T}$. Provided the underlying PDE allows spatial information to propagate, these spatial effects will appear in the time history of the sensor measurements, enabling the sensing of the entire field using only a few sensors. In vanilla SHRED, a Long Short-Term Memory (LSTM) module is used to map the sparse sensor trajectory data into a latent space, followed by a shallow decoder to reconstruct the entire spatio-temporal domain at the current time step.
SHRED enables efficient sparse sensing that is widely applicable to many scientific problems [14; 29; 45]. The advantage of SHRED comes from three aspects. First, SHRED only requires minimal sensor measurements. Under practical constraints, collecting full-state measurements for data prediction and control can be prohibitively expensive. Second, SHRED does not require grid-like data collection, which allows for generalization to more complex data structures. For example, it is easy to apply SHRED to graph data with an unknown underlying structure, such as human motion data on joints, robotic sensor data, and financial market data. Moreover, SHRED is theoretically rooted in PDE modeling methods from the perspective of separation of variables, which has the potential to offer strong theoretical guarantees such as convergence and stability.
@@ 3.1 Empowering SHRED with representation learning and physics discovery
SHRED has a natural advantage in modeling the latent governing physics due to its small model size. SHRED is based on a shallow decoder with a relatively small recurrent network structure. The relative simplicity of the model allows the latent representation to maintain many advantageous properties such as smoothness and Lipschitzness. Experimentally, we observe that the hidden state space of a SHRED model is generally very smooth. Second, SHRED does not have an explicit encoder, which avoids the potential problem of spectral bias [41]. Many reduced-order modeling methods that rely on an encoder architecture struggle to learn physics and instead focus only on modeling the low-frequency information (background) [43; 8; 36]. Building upon SHRED, we further incorporate SINDy to regularize the learned recurrence with a well-characterized and simple form of governing equation. In other words, we perform a joint discovery of a coordinate system (which transfers the high-dimensional observation into a low-dimensional representation) and the governing law (which describes how the summarized latent representation progresses forward with respect to time) of the latent space of a SHRED model. This approach is inspired by the principle in physics that, under an ideal coordinate system, physical phenomena can be described by a parsimonious dynamical model [8; 36]. When the latent representation and the governing law are well-aligned, this configuration is likely to capture the true underlying physics. This joint discovery results in a latent space that is both interpretable and physically meaningful, enabling robust and stable future prediction based on the learned dynamics.
@@ 3.2 Latent space regularization via SINDy and Koopman
As a compressive sensing procedure, there exist infinitely many equally valid solutions for the latent representation. Therefore, it is not necessary for the latent representation induced by SHRED to follow a well-structured differential equation. For instance, even if the exhibited dynamics are fundamentally linear, the latent representation may exhibit completely unexplainable dynamics, making the model challenging to interpret and extrapolate. Therefore, in SINDy-SHRED, our goal is to further constrain the latent representations to lie within the SINDy-class functional. This regularization promotes models that are fundamentally explainable by a SINDy-based ODE, allowing us to identify a parsimonious governing equation. The SINDy class of functions typically consists of a library of commonly used functions, which includes polynomials, and Fourier series. Although they may seem simple, these functions possess surprisingly strong expressive power, enabling the model to capture very complex dynamical systems.
@@ 3.2.1 SINDy as a Recurrent Neural network
Figure 2: Diagram of the RNN form of SINDy.
where $f_{\Theta}(x_{t},\Xi,\Delta t)=\Theta(x_{t})\Xi\Delta t$ is a nonlinear function. Notice that this $f_{\Theta}(\cdot)$ has exactly the same formulation as in SINDy [5]. The application of function libraries with sparsity constraints is a manner of automatic neural architecture search (NAS) [59]. Compared to all prior works [8; 18; 12], this implementation of the SINDy unit fits better in the framework of neural network training and gradient descent. We utilize trajectory data $\{{\bm{z}}_{i}\}_{i=1}^{T}$ and forward simulate the SINDy-based ODE using a trainable parameter $\Xi$. To achieve better stability and accuracy for forward integration, we use Euler integration with $k$ mini-steps (with time step $\frac{\Delta t}{k}$) to obtain ${\bm{z}}_{t+1}$. In summary, defining $h=\frac{\Delta t}{k}$, we optimize $\Xi$ with the following:
@@
To achieve $\ell_{0}$ optimization, we perform pruning with $\ell_{2}$ which approximates $\ell_{0}$ regularization under regularity conditions [58; 19; 3]. Applying SINDy unit has the following benefits: (a) The SINDy-function library contains frequently used functions in physics modeling (e.g. polynomials and Fourier series). (b) With sparse system identification, the neural network is more likely to identify governing physics, which is fundamentally important for extrapolation and long-term stability.

### v2

@@ Abstract
Modeling real-world spatio-temporal data is exceptionally difficult due to inherent high dimensionality, measurement noise, partial observations, and often expensive data collection procedures. In this paper, we present Sparse Identification of Nonlinear Dynamics with SHallow REcurrent Decoder networks (SINDy-SHRED), a method to jointly solve the sensing and model identification problems with simple implementation, efficient computation, and robust performance. SINDy-SHRED uses Gated Recurrent Units to model the temporal sequence of sparse sensor measurements along with a shallow decoder network to reconstruct the full spatio-temporal field from the latent state space. Our algorithm introduces a SINDy-based regularization for which the latent space progressively converges to a SINDy-class functional, provided the projection remains within the set. In restricting SINDy to a linear model, a Koopman-SHRED model is generated. SINDy-SHRED (i) learns a symbolic and interpretable generative model of a parsimonious and low-dimensional latent space for the complex spatio-temporal dynamics, (ii) discovers new physics models even for well-known physical systems, (iii) achieves provably robust convergence with an observed globally convex loss landscape, and (iv) achieves superior accuracy, data efficiency, and training time, all with fewer model parameters. We conduct systematic experimental studies on PDE data such as turbulent flows, real-world sensor measurements for sea surface temperature, and direct video data. The interpretable SINDy and Koopman models of latent state dynamics enable stable and accurate long-term video predictions, outperforming all current baseline deep learning models in accuracy, training time, and data requirements, including Convolutional LSTM, PredRNN, ResNet, and SimVP.
@@ 1 Introduction
Partial differential equations (PDEs) derived from first principles or qualitative behavior remain the most ubiquitous class of models used to describe physical, spatio-temporal phenomena. However, for complex dynamics it is often the case that the simplifying assumptions necessary to construct a PDE model can render it ineffectual for real data where the physics is high-dimensional, multi-scale in nature, only partially known, or where first principles models currently do not exist. In such cases, machine learning (ML) methods offer an attractive alternative for learning both the physics and coordinates (fundamental variables) necessary to quantify the observed spatiotemporal phenomena. Many recent efforts utilizing ML techniques seek to relax the computational burden for PDE simulation by learning surrogate models to forward-simulate or predict spatiotemporal systems. However, this new machine learning paradigm frequently exhibits instabilities during the training process, unstable roll-outs when modeling future state predictions, and often yields minimal computational speedups [52].
@@
The Shallow REcurrent Decoder (SHRED) network [77] is a recently introduced architecture that utilizes data from sparse sensors to reconstruct and predict the entire spatiotemporal domain. Similar to Takens’ embedding theorem, SHRED models trade spatial information at a single time point for a trajectory of sensor measurements across time. Previous work has shown SHRED can achieve excellent performance in sensing and reduced order modeling examples ranging from weather and atmospheric forecasting [77], plasma physics [41], nuclear reactors [63], and turbulent flow reconstructions [71]. Theoretically rooted in the classic PDE method of separation of variables [77, 71], the decoding-only strategy of SHRED circumvents the computation of inverse pairs, i.e. an encoder and the corresponding decoder. It has been well-known for decades that the computation of the inverse of a matrix is highly unstable and ill-posed [26, 34, 3]. By decoding only, SHRED avoids this problem and learns a single embedding without the corresponding inversion.
@@
Figure 1: (a) Illustration of the SINDy-SHRED and Koopman-SHRED architecture. SINDy-SHRED transfers the original sparse sensor signal (red) to an interpretable latent representation (purple) that falls into the SINDy-class functional. This framework can be adapted into Koopman-SHRED by restricting the library $\Theta(\cdot)$ to be linear. The shallow decoder performs a reconstruction in the state space. We obtain an interpretable linear model for the sea-surface temperature data considered (details in Sec. 4.3). (b) The loss function consists of three parts: (i) the SHRED loss controls the reconstruction accuracy of the state space, (ii) the Ensemble SINDy loss helps to model the parsimonious dynamics of the latent space, and (iii) the sparsity constraint identifies the governing equation within this optimization framework. (c) We visualize the globally convex loss landscape of SINDy-SHRED as in [43].
In this paper, we introduce Sparse Identification of Nonlinear Dynamics with SHallow REcurrent Decoder networks (SINDy-SHRED). SINDy-SHRED exploits the latent space of recurrent neural networks for sparse sensor modeling, and enforces interpretability via a SINDy-based functional class. Moreover, our theoretical analysis (see Thm. 1 and Thm. 2) rigorously demonstrates key advantages of latent space modeling with SINDy over conventional neural networks, i.e. under mild conditions it is guaranteed to converge with a bounded error. In this way, SINDy-SHRED enables a robust and sample-efficient joint discovery of governing equation and coordinate system. With the correct governing equation, SINDy-SHRED can perform an accurate long-term prediction in a learned, low-dimensional latent space, and in turn allows for long-term forecasting in the original spatiotemporal (pixel) space. In further restricting SINDy to a linear model, a Koopman approximation [9] can be constructed to produce a Koopman-SHRED architecture.
@@
SINDy-SHRED and Koopman-SHRED are lightweight models which can perform low-rank recovery with only a few active sensors for the spatio-temporal fields, which is critical for large-scale scientific data modeling and real-time control. Specifically, for $D$-dimensional fields, $D+1$ sensors are required for disambiguation of the spatio-temporal field much like localization for cellular networks [84]. Moreover, the proposed architecture does not require large amounts of data during training, thereby avoiding a common pitfall in existing ML techniques for accelerating physics simulations and enabling rapid training on a single laptop. SINDy-SHRED/Koopman-SHRED are also highly reproducible with minimal effort in hyperparameter tuning due to the globally convex optimization landscape, as illustrated in the lower right panel of Fig. 1. The recommended network structure, hyperparameters, and training setting can generalize to many different datasets  [^1] Code implementation: https://github.com/gaoliyao/sindy-shred. [^2] Run in Google Colab: https://colab.research.google.com/drive/1Xxw3P_x9a8iKZ6RPe2ZfTb8rJoWtPwTK.. In short, we demonstrate SINDy-SHRED/Koopman-SHRED to be robust and highly applicable in many modern scientific modeling problems. In what follows, we will refer generally to the proposed architecture as SINDy-SHRED, with Koopman-SHRED as a special case.
@@
We perform a wide range of studies to demonstrate the effectiveness of SINDy-SHRED. We apply the model on the sea surface temperature data, which is a complex real-world problem. We also consider data from complex simulations of atmospheric chemistry, 2D Kolmogorov flow, and isotropic turbulence, video data of flow over a cylinder, and video data of a pendulum. The ability of SINDy-SHRED to perform well on video data is an important result for so-called “GoPro physics,” whereby physics is learned directly from video recordings. With extremely small sample size and noisy environments, SINDy-SHRED achieves governing equation identification with stable long-term predictions, thus overcoming the critical limitations for long-term forecasts of spatio-temporal phenomena that include instabilities and massive computational requirements. To summarize, the contributions of our paper are the following:
@@
  • We propose SINDy-SHRED to learn a symbolic and interpretable generative model of a parsimonious and low-dimensional latent space (fundamental variables) of recurrent neural networks for complex spatio-temporal dynamics.
  • Even for the well-known physical systems considered, we discover in each case a new and hitherto unknown physics model characterizing the spatio-temporal dynamics directly from measurement or video streams. These benchmark systems include turbulent flows, videos of physical systems, and PDE simulations.
  • Under weak assumptions, we establish rigorous error bounds on the robust convergence of the SINDy-SHRED algorithm which allows us to guarantee model performance, a rarity in deep learning. Moreover, we systematically observe that the algorithm has a global convex loss landscape, preventing the need for any hyper-parameter tuning (except for the SINDy parameters of dimension and sparsity).
  • We compare SINDy-SHRED to state-of-the-art deep learning algorithms in spatio-temporal prediction, showing that the architecture achieves superior accuracy, data efficiency, training time, and long term prediction, all with fewer model parameters.
@@ 2 Related works
Traditionally, spatio-temporal physical phenomena are modeled by Partial Differential Equations (PDEs). To accelerate PDE simulations, recent efforts have leveraged neural networks to model physics. By explicitly assuming knowledge of the underlying PDE, physics-informed neural networks [60] utilize the PDE structure as a constraint for small sample learning. However, assuming the exact form of governing PDE for real data can be a strong limitation. There have been many recent works on learning and predicting PDEs directly using neural networks [38, 46, 35, 49, 47, 48]. Alternatively, variants of SINDy [64, 53, 24] offer a data-driven approach to identify PDEs from the spatial-temporal domain. High-dimensionality and data requirements can be prohibitive for many practical applications.
In parallel, efforts have been directed toward the discovery of physical laws through dimensionality reduction techniques [13, 50, 51], providing yet another perspective on the modeling of scientific data. The discovery of physics from a learned latent spaces has previously been explored by [27, 17, 23, 18, 80, 45, 58, 57, 55, 14], yet none of these methods consider a regularization on the latent space with no explicit encoder. Yu et al. proposed the idea of physics-guided learning [83] which combines physics simulations and neural network approximations. Directly modeling physics from video is also the subject of much research in the field of robotics [25, 70, 65], computer vision [81, 78, 2, 4] and computer graphics [37, liuphysgen, 79, 56], as these fields also require better physics models for simulation and control. From the deep learning side, combining the structure of differential equations into neural networks [33, 15] has been remarkably successful in a wide range of tasks. As an example, when spatial-temporal modeling is framed as a video prediction problem, He et al. found [32] that random masking can be an efficient spatio-temporal learner, and deep neural networks can provide very good predictions for the next 10 to 20 frames [67, 76, 29, 31]. Generative models have also been found to be useful for scientific data modeling [54, 69, 11, 68, 22].
@@ 3 Methods
The shallow recurrent decoder network (SHRED) is a computational technique that utilizes recurrent neural networks to estimate high-dimensional states from limited measurements [77]. The method functions by trading high-fidelity spatial information for trajectories of sparse sensor measurements at given spatial locations. Mathematically, consider a high-dimensional data series $\{{\bm{X}}_{i}\}_{i=1}^{T}\in{}^{(W\times H)\otimes T}$ that represents the evolution of a spatio-temporal dynamical system, where $W$, $H$, and $T$ denote the width, height, and total time steps of the system, respectively. In SHRED, each sensor collects data from a fixed spatial position in a discretized time domain. Denoting the subset of sensors as $\mathcal{S}$, the input data of SHRED is $\{{\bm{X}}^{\mathcal{S}}\}_{i=1}^{T}\in{}^{\text{card}(S)\otimes T}$. Provided the underlying PDE allows spatial information to propagate, these spatial effects will appear in the time history of the sensor measurements, enabling the sensing of the entire field using only a few sensors. In vanilla SHRED, a Long Short-Term Memory (LSTM) module is used to map the sparse sensor trajectory data into a latent space, followed by a shallow decoder to reconstruct the entire spatio-temporal domain at the current time step.
SHRED enables efficient sparse sensing that is widely applicable to many scientific problems [21, 41, 63, 71]. The advantage of SHRED comes from three aspects. First, SHRED only requires minimal sensor measurements. Under practical constraints, collecting full-state measurements for data prediction and control can be prohibitively expensive. Second, SHRED does not require grid-like data collection, which allows for generalization to more complex data structures. For example, it is easy to apply SHRED to graph data with an unknown underlying structure, such as human motion data on joints, robotic sensor data, and financial market data. Furthermore, SHRED is grounded in the classical technique of separation of variables [71].
@@ 3.1 Empowering SHRED with representation learning and physics discovery
SHRED has a natural advantage in modeling latent governing physics due to its small model size. SHRED is based on a shallow decoder with a relatively small recurrent network structure. The relative simplicity of the model allows the latent representation to maintain many advantageous properties such as smoothness and Lipschitzness. Experimentally, we observe that the hidden state space of a SHRED model is generally very smooth. Second, SHRED does not have an explicit encoder, which avoids the potential problem of spectral bias [59]. Many reduced-order modeling methods that rely on an encoder architecture struggle to learn physics and instead focus only on modeling the low-frequency information (background) [61, 13, 51]. Building upon SHRED, we further incorporate SINDy to regularize the learned recurrence with a well-characterized and simple form of governing equation. This approach is inspired by the principle in physics that, under an ideal coordinate system, physical phenomena can be described by a parsimonious dynamics model [13, 51]. When the latent representation and the governing law are well-aligned, this configuration is likely to capture the true underlying physics. This joint discovery results in a latent space that is both interpretable and physically meaningful, enabling robust and stable future prediction based on the learned dynamics.
@@ 3.2 Latent space regularization via SINDy and Koopman
As a compressive sensing procedure, there exist infinitely many equally valid solutions for the latent representation. Therefore, it is not necessary for the latent representation induced by SHRED to follow a well-structured differential equation. For instance, even if the exhibited dynamics are fundamentally linear, the latent representation may exhibit completely unexplainable dynamics, making the model challenging to interpret and extrapolate. Therefore, in SINDy-SHRED, our goal is to further constrain the latent representations to lie within the SINDy-class functional. This regularization promotes models that are fundamentally explainable by a SINDy-based ODE, allowing us to identify a parsimonious governing equation. The SINDy class of functions typically consists of a library of commonly used functions, which includes polynomials and Fourier series. Although they may seem simple, these functions possess surprisingly strong expressive power, enabling the model to capture very complex dynamical systems.
@@ 3.2.1 SINDy as a Recurrent Neural network
where $f_{\Theta}(x_{t},\Xi,\Delta t)=\Theta(x_{t})\Xi\Delta t$ is a nonlinear function. Notice that this $f_{\Theta}(\cdot)$ has exactly the same formulation as in SINDy [10]. We demonstrate a visual form of this design in the Appendix. The application of function libraries with sparsity constraints is a manner of automatic neural architecture search (NAS) [87]. Compared to all prior works [13, 27, 18], this implementation of the SINDy unit fits better in the framework of neural network training and gradient descent. We utilize trajectory data $\{{\bm{z}}_{i}\}_{i=1}^{T}$ and forward simulate the SINDy-based ODE using a trainable parameter $\Xi$. To achieve better stability and accuracy for forward integration, we use Euler integration with $k$ mini-steps (with time step $\frac{\Delta t}{k}$) to obtain ${\bm{z}}_{t+1}$. In summary, defining $h=\frac{\Delta t}{k}$, we optimize $\Xi$ with the following:
@@
To achieve $\ell_{0}$ optimization, we perform pruning with $\ell_{2}$ regularization which is known to approximate $\ell_{0}$ regularization under regularity conditions [86, 28, 8]. Applying SINDy unit has the following benefits: (a) The SINDy-function library contains frequently used functions in physics modeling (e.g. polynomials and Fourier series). (b) With sparse system identification, the neural network is more likely to identify governing physics, which is fundamentally important for extrapolation and long-term stability.